SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (July 18, 2006)

Attendees: Rainer Bartoldus, Selina Li, Rahmat Rahmat, Dong Su, Dan Walker

Operations

  • On July 14 (last Friday), the L1 information failed to display in EPICS panels because the FCPM stopped reading it's monitoring data. TRG was paged for this, but Selina suggested they paged ODF instead since the L3 rate seemed to indicate that the trigger was working fine. According to Matt Meaver (ODF), this happened once before to another fast control partition module (FCPM). Restarting ORC solved the problem.
  • Yesterday (July 17), PEP was down to replace some bellows that were suspected to be the reasons for beam aborts.
  • Selina will give a monthly TRG status report after tomorrow's detector operation meeting at IR2. A link to the talk is provided here.

EMT

  • Dan found out that the EMT deadtime information is in the database since the 10th of October, 2005. Dan showed a few plots for the EMT deadtime for different run period. He will verify with VMON to see if the FEX suppression actually helps.

DCT

  • The TSF-Y fiber connection was removed from the splitter and now the fibers are directly connected from DCH to TSF-i during the PEP downtime on July 17 (Monday). See HN posted by Su Dong. Cosmic runs were taken before the beams came back to make sure no damage was made during this process. The optical power for Y doubled after the removal of the splitter. Selina showed two screen shots for comparing before and after the removal.
  • Su Dong found out that the L1Ntuple failed to get the TSFLUT when the debug TSF data are involved. We suspected that the LUT file might have been moved. Rainer reminded us that the TSFLUT can be retrieved from database.

Online

  • Release 18.7.0 has been validated except for skimming. This release should install at IR2 soon so some of the changes we made in L3 and L1TFMon will take effect.
  • Rainer will post a HN to propose a change in the scalar group so we could have time history plots for 60 minutes rather than only 45 minutes of a run.

Data Quality and Monitoring

  • Rahmat showed an example of strip chart for the lumi ratio of gamma gamma and mu-pair. Note that the numbers are from beta lumi. There will be a total of 4 strip charts for lumi ratios. Rahmat is aiming to put all 4 lumi ratio in one plot into OPR release for production by next week. Al suggested that maybe we can show these strip charts at the validation meeting.

Trigger Configuration

  • Rainer posted some new changes to the trigger statistics package. See HN posted. He is writing up some documentation for this trigger statistics tool. After getting inputs from analysis users, we will work on a better configuration to test before the end of run 5b.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.